# Dexonomy: Synthesizing All Dexterous Grasp Types in a Grasp Taxonomy

### v1

@@ I Introduction
In this work, we propose a novel pipeline based on sampling and optimization to address these challenges. As shown in Figure LABEL:fig:teaser, our algorithm can efficiently synthesize high-quality dexterous grasps for any grasp type, object, and articulated hand, requiring only one human-annotated grasp template per hand and grasp type. Our synthesized grasps achieve rich hand-object contact (e.g., $>10$ hand links within 2 mm of the object for power grasps), guarantee penetration-free poses via collision mesh verification, and satisfy force closure under six-axis testing in MuJoCo [42] — all with shared hyperparameters across grasp types, objects, and hands.
@@ III Preliminary of Grasp Quality Metric
$\mathcal{F}_{i}$ =$=\left\{\mathbf{x}_{i}\in\mathbb{R}^{3}~{}|~{}0\leq x_{i,1}\leq 1,x_{i,2}^{2}+x_{i,3}^{2}\leq\mu^{2}x_{i,1}^{2}\right\}$ (1)
@@ IV-C Simulation-based Local Refinement of Hand Pose
[…]
@@ V Dexonomy Dataset
Using our proposed grasp synthesis pipeline, we construct a large-scale dataset for Shadow hand covering 31 grasp types from the GRASP taxonomy [20]. This dataset is designed to support research on data-driven methods for type-aware grasp synthesis. Two grasp types in the taxonomy, Distal Type (#$\#19$) and Tripod Variation (#$\#21$), are excluded due to their specificity to object categories, namely scissors and chopsticks, respectively.
@@
[…]
@@
[…]
@@ VII-A Evaluation Metrics
[…]
@@
[…]
@@ VII-C1 Comparison with functional grasp transfer baselines
TABLE IV: Statistics of Grasp Synthesis for the GRASP Taxonomy. The success rate is lower than fingertip grasps because many flexible grasp types are suitable only for specific objects, e.g., Lateral (#$\#16$) grasps for flat objects.
@@ VII-C2 Statistics analysis of our pipeline
The overall success rate is considerably lower than that of fingertip grasp synthesis, as many flexible grasp types are designed for specific object shapes. For instance, the Lateral (#$\#16$) grasp is only used for flat and small objects. Among different grasp types, precision grasps exhibit the highest success rate under normal test conditions (i.e., with friction coefficients of 0.6 and 0.02), since these grasps typically involve only the fingertips and suit more objects. However, the success rate of precision grasps drops more rapidly than that of power grasps when the friction coefficients are reduced to 0.3 and 0.002 (i.e., the hard test conditions), indicating that power grasps offer higher stability due to more contact with the object. Additionally, the overall diversity of grasps is better than previous work reported in Table II, owing to the inclusion of many distinct grasp types.
@@ VII-D1 Modular ablation
Table V shows the impact of each module on fingertip grasp synthesis for the Allegro hand. For the global alignment stage, optimization provides a slight improvement, while post-filtering plays a more significant role. We also take a deeper look at the post-filtering of this stage, and find that a significant portion ($>50$k out of $80$k) is filtered out due to a high loss (Eq. 7). In contrast, collision and grasp quality filters only remove fewer than 10k grasps, indicating that using them as new losses for optimization is not effective. In the local refinement stage, optimization is crucial for improving grasp quality, while post-filtering primarily ensures penetration-free grasps and negatively affects the success rate. Additionally, constructing new grasp templates significantly enhances the synthesis success rate.
[…]
@@ VII-D2 Control strategy ablation
In Table VI, we evaluate the impact of different control strategies during simulation testing. The Grasp only method computes the pre-grasp and squeeze joint angles by scaling the grasp joint angles with fixed factors: 0.9×0.9$\times$0.9 × for the pre-grasp and $\times1$ for the squeeze pose. The w/ pre-grasp method calculates the squeeze joint angles as $squeeze=2\times grasp-pregrasp$, following BODex [7]. Note that these two strategies only calculate the joint angles, while the 6DoF hand root pose is the same as the one of the grasp pose. During execution, the hand transitions sequentially from the pre-grasp to the grasp pose, and then to the squeeze pose.
@@ VII-E Learning-based Grasp Synthesis from Partial Observation
In this section, we compare the influence of both the grasping dataset and the learning method in simulation. The 10.7k objects in our Dexonomy dataset are randomly split into training and test sets with a 4:1 ratio. While the object scales used for training vary, we fix the scales during testing, using the same six scale levels as described in Section VII-B. To ensure a fair comparison, we also regenerate a dataset for BODex using our objects and scales, resulting in 0.7M valid grasps. Ours-type1 includes only the Large Diameter (#$\#1$) grasp type from the Dexonomy dataset and contains 0.4M data points, while Ours-all uses the full 9.5M dataset. For the type-conditional model, we additionally train a classifier to select the best grasp type based on each object’s point cloud; more details are provided in Appendix -D. For each object, 100 candidate grasps are predicted and ranked by their associated probabilities, with the top 10 selected as the final outputs.
@@ VII-F Real-World Experiment
As shown in Figure 7, our model can correctly generate physically plausible grasps for the specified types and achieves an overall success rate of $82.3\%$. The most common failure mode is the unsuitable grasp types for the object, especially for Lateral (#$\#16$). Moreover, some grasp types, such as Tip Pinch (#$\#24$), are not very robust and easily fail due to real-world noise and prediction errors. More details are provided in Appendix -E.
@@ Acknowledgments
We thank Danshi Li for assisting with the integration of the Inspire Hand and the Unitree G1 Hand.

### v2

@@ I Introduction
In this work, we propose a novel pipeline based on sampling and optimization to address these challenges. As shown in Figure LABEL:fig:teaser, our algorithm can efficiently synthesize high-quality dexterous grasps for any grasp type, object, and articulated hand, requiring only one human-annotated grasp template per hand and grasp type. Our synthesized grasps achieve rich hand-object contact (e.g., >10>10 hand links within 2 mm of the object for power grasps), guarantee penetration-free poses via collision mesh verification, and satisfy force closure under six-axis testing in MuJoCo [42] — all with shared hyperparameters across grasp types, objects, and hands.
@@ III Preliminary of Grasp Quality Metric
$\mathcal{F}_{i}$ $=\left\{\mathbf{x}_{i}\in\mathbb{R}^{3}~|~0\leq x_{i,1}\leq 1,x_{i,2}^{2}+x_{i,3}^{2}\leq\mu^{2}x_{i,1}^{2}\right\}$ (1)
@@ IV-C Simulation-based Local Refinement of Hand Pose
To ensure the hand remains strictly penetration-free with respect to the object, we apply a $1\text{mm}$ contact margin in MuJoCo—meaning the hand is repelled if it comes within $1\text{mm}$ of the object surface. Since MuJoCo reliably resolves penetration at the millimeter scale, the resulting contact distance between the hand and the object typically falls within the range of $0,2\text{mm}]$. This behavior motivates our use of a $2\text{mm}$ contact tolerance: a hand link is considered to be in contact if its distance to the object is less than $2\text{mm}$.
@@ V Dexonomy Dataset
Using our proposed grasp synthesis pipeline, we construct a large-scale dataset for Shadow hand covering 31 grasp types from the GRASP taxonomy [20]. This dataset is designed to support research on data-driven methods for type-aware grasp synthesis. Two grasp types in the taxonomy, Distal Type (#19\#19) and Tripod Variation (#21\#21), are excluded due to their specificity to object categories, namely scissors and chopsticks, respectively.
@@
  • Pre-grasp pose for collision-free motion planning, generated after the grasp pose by enforcing a $2\text{cm}$ contact margin in MuJoCo—pushing the hand away if it is within $2\text{cm}$ of the object.
@@
GSR ($\%$) $\uparrow$ OSR ($\%$) $\uparrow$ S ($s^{-1}$) $\uparrow$ CLN $\uparrow$ CDC ($mm$) $\downarrow$ PD ($mm$) $\downarrow$ SPD ($mm$) $\downarrow$ D ($\%$) $\downarrow$
@@ VII-A Evaluation Metrics
Grasp Success Rate (GSR) (unit: $\%$): The percentage of successful grasps relative to the attempt number. For our method, one attempt is defined as one valid result output by the global alignment stage. A grasp succeeds only if it resists six external forces in MuJoCo and does not have severe penetrations (>1>1 cm), since the penetration may cause simulation failure and prevent the object from moving. The object mass is 100g, and the success criteria for the object pose are $5\text{cm}$ and $15^{\circ}$.
Object Success Rate (OSR) (unit: $\%$): The percentage of objects that have at least one successful grasp. If the object scales are fixed, as in Sections VII-B and VII-E, different scales of the same object are treated as separate objects.
@@
Diversity (D) (unit: $\%$): The proportion of total variance explained by the first principal component in PCA, computed as the ratio of the first eigenvalue to the sum of all eigenvalues. PCA is performed on data points that include grasp translation $\mathbf{T}_{g}$, rotation $\mathbf{R}_{g}$ (in the axis-angle representation), and joint angles $\mathbf{q}_{g}$.
@@ VII-C1 Comparison with functional grasp transfer baselines
TABLE IV: Statistics of Grasp Synthesis for the GRASP Taxonomy. The success rate is lower than fingertip grasps because many flexible grasp types are suitable only for specific objects, e.g., Lateral (#16\#16) grasps for flat objects.
@@ VII-C2 Statistics analysis of our pipeline
The overall success rate is considerably lower than that of fingertip grasp synthesis, as many flexible grasp types are designed for specific object shapes. For instance, the Lateral (#16\#16) grasp is only used for flat and small objects. Among different grasp types, precision grasps exhibit the highest success rate under normal test conditions (i.e., with friction coefficients of 0.6 and 0.02), since these grasps typically involve only the fingertips and suit more objects. However, the success rate of precision grasps drops more rapidly than that of power grasps when the friction coefficients are reduced to 0.3 and 0.002 (i.e., the hard test conditions), indicating that power grasps offer higher stability due to more contact with the object. Additionally, the overall diversity of grasps is better than previous work reported in Table II, owing to the inclusion of many distinct grasp types.
@@ VII-D1 Modular ablation
Table V shows the impact of each module on fingertip grasp synthesis for the Allegro hand. For the global alignment stage, optimization provides a slight improvement, while post-filtering plays a more significant role. We also take a deeper look at the post-filtering of this stage, and find that a significant portion (>50>50k out of 8080k) is filtered out due to a high loss (Eq. 7). In contrast, collision and grasp quality filters only remove fewer than 10k grasps, indicating that using them as new losses for optimization is not effective. In the local refinement stage, optimization is crucial for improving grasp quality, while post-filtering primarily ensures penetration-free grasps and negatively affects the success rate. Additionally, constructing new grasp templates significantly enhances the synthesis success rate.
GSR ($\%$) $\uparrow$ Shadow Allegro
@@ VII-D2 Control strategy ablation
In Table VI, we evaluate the impact of different control strategies during simulation testing. The Grasp only method computes the pre-grasp and squeeze joint angles by scaling the grasp joint angles with fixed factors: $0.9\times$ for the pre-grasp and $1.1\times$ for the squeeze pose. The w/ pre-grasp method calculates the squeeze joint angles as $squeeze=2\times grasp-pregrasp$, following BODex [7]. Note that these two strategies only calculate the joint angles, while the 6DoF hand root pose is the same as the one of the grasp pose. During execution, the hand transitions sequentially from the pre-grasp to the grasp pose, and then to the squeeze pose.
@@ VII-E Learning-based Grasp Synthesis from Partial Observation
In this section, we compare the influence of both the grasping dataset and the learning method in simulation. The 10.7k objects in our Dexonomy dataset are randomly split into training and test sets with a 4:1 ratio. While the object scales used for training vary, we fix the scales during testing, using the same six scale levels as described in Section VII-B. To ensure a fair comparison, we also regenerate a dataset for BODex using our objects and scales, resulting in 0.7M valid grasps. Ours-type1 includes only the Large Diameter (#1\#1) grasp type from the Dexonomy dataset and contains 0.4M data points, while Ours-all uses the full 9.5M dataset. For the type-conditional model, we additionally train a classifier to select the best grasp type based on each object’s point cloud; more details are provided in Appendix -D. For each object, 100 candidate grasps are predicted and ranked by their associated probabilities, with the top 10 selected as the final outputs.
@@ VII-F Real-World Experiment
As shown in Figure 7, our model can correctly generate physically plausible grasps for the specified types and achieves an overall success rate of $82.3\%$. The most common failure mode is the unsuitable grasp types for the object, especially for Lateral (#16\#16). Moreover, some grasp types, such as Tip Pinch (#24\#24), are not very robust and easily fail due to real-world noise and prediction errors. More details are provided in Appendix -E.
@@ Acknowledgments
This work was supported by Beijing Natural Science Foundation (Grant No.QY24042). We also thank Danshi Li for assisting with the integration of the Inspire Hand and the Unitree G1 Hand.